Seismicity during this period was low, with only one volcano-tectonic earthquake being recorded by the network. This event was located at shallow depth beneath the crater. Small long-period events have been recorded only on Gages station throughout the period. There was one period of continuous, low amplitude broadband tremor between 22:00 and midnight on 16 January and there have been several other short periods during the day today. An observed rock-fall event at 13:10 today correlated well with a signal recorded at Gages and Long Ground; several similar signals have been recorded throughout the day at the Gages station which are interpreted as other rock fall events.
EDM measurements were made today on the Long Ground - White's Yard - Castle Peak triangle. Initial analysis of the results suggests no change in any of the line lengths since the last occupation two days ago.
Visual observations were made this afternoon from the helicopter and from the ground at Harris and Long Ground. A new spine has grown quickly over the past 36 to 48 hours close to the remnants of the one that collapsed a few days ago. An area of very active rock falling has developed to the north of this spine, and it may collapse within the next day or so. There were several areas of vigorous steaming on the new dome and the entire structure appears still to be swelling slowly.
Richard Robertson, Stefan Sol and Desmond Seupersad arrived today from SRU in Trinidad to replace Dr William Ambeh and Chandradath Ramsingh who will be leaving the MVO team at the end of the week.